# CounselBench: A Large-Scale Expert Evaluation and Adversarial Benchmark of Large Language Models in Mental Health Counseling

### v1

@@ Abstract
Large language models (LLMs) are increasingly proposed for use in mental health support, yet their behavior in realistic counseling scenarios remains largely untested. We introduce CounselBench, a large-scale benchmark developed with 100 mental health professionals to evaluate and stress-test LLMs in single-turn counseling. The first component, CounselBench-EVAL, contains 2,000 expert evaluations of responses from GPT-4, LLaMA 3, Gemini, and online human therapists to real patient questions. Each response is rated along six clinically grounded dimensions, with written rationales and span-level annotations. We find that LLMs often outperform online human therapists in perceived quality, but experts frequently flag their outputs for safety concerns such as unauthorized medical advice. Follow-up experiments show that LLM judges consistently overrate model responses and overlook safety issues identified by human experts. To probe failure modes more directly, we construct CounselBench-Adv, an adversarial dataset of 120 expert-authored counseling questions designed to trigger specific model issues. Evaluation across 2,880 responses from eight LLMs reveals consistent, model-specific failure patterns. Together, CounselBench establishes a clinically grounded framework for benchmarking and improving LLM behavior in high-stakes mental health settings.
@@
Disclaimer: this study does not endorse or recommend the use of LLMs in therapeutic practice. Our goal is to evaluate and characterize LLM behavior in counseling interactions for research only.
@@ 1 Introduction
Professional counseling is a critical pillar of mental healthcare, yet remains inaccessible to many due to structural barriers such as cost, limited provider availability, transportation challenges, and stigma DockeryEtAl2015 ; GargEtAl2022a ; Gary2005 ; TristianaEtAl2018 . Digital technologies have improved access through tools such as smartphone-based apps NearySchueller2018 , tele-therapy (telemedicine_mental_health, ) and online counseling services MarcelleEtAl2019 ; JoEtAl2023 , which allows users to seek support anonymously, affordably, and on demand NearySchueller2018 ; SagarPattanayak2015 ; ZhangEtAl2015 . However, even with these innovations, many individuals still lack adequate support due to difficualty finding a relatable therapist, distrust of the mental health systems NeukrugEtAl2017 ; JardineEtAl2024a , or a preference for peer-driven advice RepperandCarter2011 ; Gillard2019 . As a result, people often turn to general-purposed platforms like Reddit (mental_disorder_reddit, ), where the quality of advice is inconsistent and unregulated.
Recent advances in natural language processing (NLP) have introduced a new player in this space: large language models (LLMs). These models undergo large-scale pretraining on diverse data and have the ability to understand context and provide immediate answers, which presents a compelling opportunity to assist with mental health support chen2023soulchatimprovingllmsempathy ; Stade_Stirman_Ungar_Boland_Schwartz_Yaden_Sedoc_DeRubeis_Willer_Eichstaedt_2024_llm_change_behavioral_healthcare ; hua2024largelanguagemodelsmental . Individuals may engage directly with LLMs, such as ChatGPT, for psychological support (LLM_as_mh_resources, ) or through derivative applications specifically designed for mental healthcare (opportunities_risks_llm_mental_health, ).
Despite this rapid deployment, little is known about how LLMs behave in real-world counseling contexts. Recent incidents and critiques have raised concerns about inappropriate advice, lack of safeguards, and misalignment with clinical standards character_ai_lawsuit ; Duffy_2024 . Most existing medical NLP benchmarks focus on structured tasks like diagnosis or factual question answering galatzerlevy2023capabilitylargelanguagemodels ; nori2023capabilitiesgpt4medicalchallenge ; Xu_2024_mental_llm . However, effective counseling requires a different set of capabilities, such as emotional nuance, contextual sensitivity, and open-ended reasoning. Prior work on mental health dialogue has largely focused on simulated multi-turn conversations with small-scale human annotation wang2024clientcenteredassessmentllmtherapists ; iftikhar2024therapynlptaskpsychologists ; chaszczewicz2024multilevelfeedbackgenerationlarge ; wang2024patientpsiusinglargelanguage , leaving key gaps in evaluating how LLMs respond to authentic user concerns in a realistic setting.
In this work, we focus on a common and under-explored setting: single-turn counseling, where an individual asks a counseling question and receives a standalone response. This format reflects real-world use cases on online forums, counseling Q&A platforms, and chat-based interventions JAMA-patient-clinician-interaction2024 , and is particularly relevant for Just-in-Time Adaptive Interventions (JITAIs) just_in_time_adaptive_intervention . Because there is no opportunity for follow-up or repair, quality and safety in this setting are especially important.
We aim to answer three key questions:
@@
    How do LLM-generated counseling responses compare to those from online human therapists in perceived quality and safety?
@@
    Can LLMs reliably evaluate the quality of counseling responses?
@@
    What kinds of failure modes do LLMs exhibit in counseling settings, and can those failure modes be systematically elicited?
@@ 2 Overview of CounselBench
To address these questions, we construct CounselBench in collaboration with 100 mental health professionals. COUNSELBENCH is a two-part benchmark for evaluating LLMs in realistic, single-turn mental health counseling scenarios (Figure 1). It includes:
  • COUNSELBENCH-Eval: a benchmark of 2,000 expert evaluations of LLMs and online human therapist responses to real counseling questions, with written rationales and span-level annotations.
@@
Figure 1: Overview of CounselBench benchmark. COUNSELBENCH-EVAL (left) includes expert evaluation of LLMs and human responses to real counseling questions. COUNSELBENCH-ADV (right) includes adversarial questions authored by clinicians to target identified LLM failure modes. See Appendix B for distinct license/degree types and specialization areas.
Research Pipeline
We began by extracting 100 real counseling questions from a publicly available dataset scraped from CounselChat bertagnolli2020counselchat , covering 20 common mental health topics (§3.1). Each question was answered by GPT-4, Gemini, LLaMA-3, and an online human therapist (§3.2). To evaluate these responses, we defined six evidence-based dimensions grounded in counseling literature: overall quality, empathy, specificity, factual consistency, medical advice, and toxicity (§3.3).
We then recruited 100 mental health professionals, covering 43 categories of counseling expertise, to evaluate the quality and safety of each response (§3.4). Besides rating the messages using numeric scores, they were asked to provide span-level annotation and in-depth elaboration of their rating rationales, offering rich, contextual insights into how professionals formulate counseling messages deliberately. We then analyzed these expert evaluations to compare counseling quality across models and human responses (§3.5). The evaluations of human and model responses form the COUNSELBENCH-EVAL dataset. In parallel, we prompted eight LLMs to rate the same responses using the same criteria (§3.6).
To stress-test model safety further, we asked 10 experts to author 120 adversarial questions that are likely to surface the common LLM failure modes identified in COUNSELBENCH-EVAL. These questions were used to test eight LLMs under targeted issues and form the COUNSELBENCH-ADV dataset (§4). Together, CounselBench enables both broad benchmarking and fine-grained analysis of LLM performance, including comparison with human and automated evaluations.
3 CounselBench-EVAL: Benchmarking Counseling Quality with 100 Mental Health Professionals and 8 LLMs
@@
We develop COUNSELBENCH-EVAL to systematically assess LLM behavior in realistic, single-turn counseling scenarios. This section describes the construction pipeline: question sourcing (§3.1), LLM response generation (§3.2), evaluation rubric design (§3.3), human annotation protocol (§3.4), results from expert scoring and analysis (§3.5), and experiments with LLM-based evaluators (§3.6).
3.1 Real-World Counseling Questions from Patients
Our study collects real-world counseling inquiries and responses between patients and clinicians from CounselChat, an online counseling platform where licensed therapists publicly respond to anonymous patient inquiries. We build on a version of the dataset compiled by bertagnolli2020counselchat , covering patient-clinician interactions posted before 2022. All online human responses were written by licensed human therapists before the widespread adoption of powerful LLMs like ChatGPT, which was released in late 2022.
@@ 3.1 Real-World Counseling Questions from Patients
To ensure topic diversity and response quality, we curated a diverse set of 100 counseling questions, with five questions in each of 20 topics: (1) depression, (2) relationships, (3) anxiety, (4) family conflict, (5) parenting, (6) self esteem, (7) relationship dissolution, (8) behavioral change, (9) anger management, (10) trauma, (11) marriage, (12) domestic violence, (13) grief and loss, (14) social relationships, (15) workplace relationships, (16) legal regulatory, (17) substance abuse, (18) counseling fundamentals, (19) eating disorders, (20) professional ethics. Full selection and filtering criteria are detailed in Appendix LABEL:appendix_qa_selection.
For each question, we selected the top upvoted therapist response as the "online human therapist" entry in our dataset. While upvotes are not a clinical quality measure, they offer a practical, publicly observable signal of relevance and perceived helpfulness in peer-supported mental health forums.
@@ 3.2 Generating Counseling Responses from LLMs
To ensure fair comparison and reduce annotator fatigue, we limited our evaluation to three LLMs per question in COUNSELBENCH-EVAL. Each mental health professional blindly annotated all four responses to the same question (one human-written and three LLM-generated), allowing consistent side-by-side assessment while keeping the task within a manageable cognitive load.
We selected three representative LLMs for evaluation: (1) GPT-4-0613 (via OpenAI API), a widely adopted commercial model, (2) LLaMA-3.3-70B-Instruct (accessed through Hugging Face), a leading open-source model, (3) Gemini-1.5-Pro (accessed via Google’s Generative AI SDK), which powers responses in Google Search and Assistant. Together, they represent major deployment paradigms: commercial APIs, open-source checkpoints, and search-integrated platforms. Each model is prompted with the original patient question and asked to generate a single-turn counseling response. Full prompt templates and generation details are provided in Appendix C.1.
@@ 3.3 Evaluation Rubric and Paradigms
To assess the quality and safety of counseling responses, we developed a multi-dimensional evaluation rubric based on clinical psychology literature and expert consultation. All metrics were rated using 5-point Likert scales ($1=$ the most negative; $5=$ the most positive) unless specified otherwise. Appendix A shows a screenshot of the rating questions and scales for evaluation.
@@
Empathy measures whether the response demonstrates emotional attunement, compassion, or validation. Empathy is a central component of therapeutic modalities (e.g., person-centered therapy JaniEtAl2012 ), and is strongly linked to therapeutic alliance, a key predictor of positive therapy outcomes AckermanHilsenroth2003 .
Specificity evaluates how well the response is tailored to the user’s particular context, rather than offering generic or overly broad advice. Personalized replies help users feel heard and understood (ElkinandMcKay2014, ; StilesHorvath2017, ; KramerStiles2015, ), and contextual relevance is critical to effective support (HitchcockEtAl2024, ; StilesHorvath2017, ; KramerStiles2015, ).
Medical Advice flags whether the response includes therapeutic or diagnostic guidance that should only be provided by licensed professionals. Such advice typically requires individualized assessment and clinical expertise (FrankEtAl2020, ; LyonEtAl2011, ), and its timing and appropriateness must be carefully managed by trained practitioners (HillEtAl2023, ; PrassandKivlighanJr.2021, ). This dimension is treated as binary (Yes/No) with an additional “I am not sure" option, and higher rates of flagged content indicate safety concerns.
Factual Consistency checks whether the response aligns with accepted commonsense or clinical knowledge and avoids inaccurate or unsupported claims. Users often struggle to discern factual accuracy in online mental health content (Morahan-MartinAnderson2000, ). LLMs are known to hallucinate false information (SunEtAl2025, ), and even human counselors can produce factual errors in communication (Lobel2024, ). This dimension is rated on a 4-point scale, also with an additional “I’m not sure” option to accommodate cases where the rater cannot confidently assess it.
Toxicity assesses the presence of language that is potentially harmful, stigmatizing, dismissive, or ethically problematic. Individuals seeking mental health support are often in vulnerable states and may be particularly susceptible to damaging content (WebbEtAl2008, ; MitchellEtAl2017, ).
Rating Rationale & Text Extraction
To capture the context and reasoning behind human evaluations of counseling messages, we asked mental health professionals (§3.4) to provide open-text explanations for their ratings on Overall Quality, Medical Advice, Factual Consistency, and Toxicity. Additionally, we asked them to identify and extract specific portions of each counseling response that supported their ratings. Two researchers on the team checked and verified the quality of text responses after data collection. In general, human annotators provided detailed written feedback, with an average length of more than $600$ words per survey combining all rationales.

### v2

# CounselBench: A Large-Scale Expert Evaluation and Adversarial Benchmarking of Large Language Models in Mental Health Question Answering

@@ Abstract
Medical question answering (QA) benchmarks often focus on multiple-choice or fact-based tasks, leaving open-ended answers to real patient questions underexplored. This gap is particularly critical in mental health, where patient questions often mix symptoms, treatment concerns, and emotional needs, requiring answers that balance clinical caution with contextual sensitivity. We present CounselBench, a large-scale benchmark developed with 100 mental health professionals to evaluate and stress-test large language models (LLMs) in realistic help-seeking scenarios. The first component, CounselBench-EVAL, contains 2,000 expert evaluations of answers from GPT-4, LLaMA 3, Gemini, and human therapists on patient questions from the public forum CounselChat. Each answer is rated across six clinically grounded dimensions, with span-level annotations and written rationales. Expert evaluations show that while LLMs achieve high scores on several dimensions, they also exhibit recurring issues, including unconstructive feedback, overgeneralization, and limited personalization or relevance. Responses were frequently flagged for safety risks, most notably unauthorized medical advice. Follow-up experiments show that LLM judges systematically overrate model responses and overlook safety concerns identified by human experts. To probe failure modes more directly, we construct CounselBench-Adv, an adversarial dataset of 120 expert-authored mental health questions designed to trigger specific model issues. Evaluation of 3,240 responses from nine LLMs reveals consistent, model-specific failure patterns. Together, CounselBench establishes a clinically grounded framework for benchmarking LLMs in mental health QA.
@@
Disclaimer: this study is conducted for research purposes and is not intended for clinical use or as a substitute for professional medical advice.
@@ 1 Introduction
Large language models (LLMs) have shown impressive progress in general-purpose question answering (QA), and there is growing interest in their use for healthcare applications (Thirunavukarasu et al., 2023). However, deployment in high-stakes settings requires rigorous evaluation of both capability and risk (Ali et al., 2023). Existing medical QA benchmarks such as MedQA (Jin et al., 2020) and MedMCQA (Pal et al., 2022) largely emphasize multiple-choice or fact-based tasks (Zhang et al., 2025). While these benchmarks measure factual recall, they fail to capture how models perform on the open-ended questions that patients actually ask. Open-ended QA presents greater challenges: questions are ambiguous, responses cannot be reduced to a single correct answer, and evaluation requires criteria that extend beyond factual accuracy.
These challenges are particularly critical in mental health QA. Here, patient queries are expressed as free-text descriptions of symptoms, treatment concerns, and emotional needs, and helpful responses must balance empathy, actionable guidance, and professional boundaries (Elliott et al., 2011; Na, 2024; Meadi et al., 2025). Unlike other medical domains, where diagnostic tests or structured knowledge often constrain the answer space, mental health QAs are inherently subjective and context-dependent (Association, 2013; Hormazábal-Salgado et al., 2024). Moreover, mental health QA has direct real-world significance: platforms such as CounselChat (Bertagnolli, 2020), peer-support forums (Pendse et al., 2019), EHR-based messaging, and digital mental health services like NOCD (NOCD, 2018) provide brief, asynchronous answers at scale. For many individuals, these platforms are a primary source of support (Naslund et al., 2016; Merchant et al., 2022). In such settings, even small errors, such as speculative claims, unsafe recommendations, or dismissive tone, can have immediate negative consequences. Evaluating LLMs on authentic, open-ended mental health QA is therefore both technically and socially urgent.
While recent work has begun to explore mental health QA, current resources remain limited. Many benchmarks continue to rely on multiple-choice proxies, which bypass the ambiguity and contextual demands of free-text responses (Racha et al., 2025). Others use small expert panels or LLM-as-judge protocols (Sun et al., 2021; Szymanski et al., 2024; Racha et al., 2025), approaches that provide efficiency but often miss clinically salient failures. These limitations highlight the need for a large-scale, clinician-grounded benchmark for open-ended mental health QA, as well as resources designed to stress-test LLMs and surface failure modes that matter in practice.
We introduce CounselBench to address this gap. Our contributions are threefold:
@@
    Clinically grounded evaluation dimensions: we define six criteria for open-ended mental health QA, capturing both quality and safety.
@@
    CounselBench-Eval: we recruit 100 licensed mental health professionals to provide 2,000 evaluations of answers to real patient questions, with span-level annotations and written rationales, making it among the largest expert-annotated datasets for mental health QA.
@@
    CounselBench-Adv: we construct an adversarial dataset of 120 expert-authored mental health questions designed to elicit systematic failure modes, providing targeted signals for model improvement.
Together, these contributions move evaluation beyond factual correctness toward clinically meaningful behavior, establishing a practitioner-anchored benchmark for LLMs in mental health QA.
2 Related Work
LLMs in Medical QA Benchmarks. Early medical QA benchmarks emphasized closed-form accuracy, such as multiple-choice exams (Pal et al., 2022) or retrieval-based tasks (Zhu et al., 2020). Recent benchmarks broaden evaluation with multi-axis rubrics. For example, MultiMedQA (Singhal et al., 2022) introduces physician-designed criteria such as factuality and bias. HealthBench (Arora et al., 2025) further scales to tens of thousands of physician-curated items in general medical domains. These works advance evaluation beyond accuracy but remain focused on structured medical knowledge. In contrast, our work targets authentic mental health QA, where
Evaluating Open-Ended Medical QA. Evaluating free-form medical QA remains difficult. Automated metrics (e.g., BLEU) scale easily but poorly capture qualities like specificity or safety (Croxford et al., 2024). Some studies incorporate expert review, but often with small panels due to cost (Hosseini et al., 2024; Kim et al., 2024; Manes et al., 2024). LLM-as-judge approaches are emerging (Szymanski et al., 2024), yet their reliability in clinical domains requires human validation. Many works also use non-experts or heuristics (Turcan & McKeown, 2019), risking inconsistent results.
Benchmarks for Mental Health QA. Existing mental health QA datasets mostly focus on exam-style knowledge tests with objective answer keys (Racha et al., 2025; Nguyen et al., 2025). Others adopt psychological rubrics (Li et al., 2024; Chiu et al., 2024) but rarely involve clinicians in design, often relying on crowdworkers or heuristics (Young et al., 2024). Our evaluation protocol is co-designed with clinical psychologists and scaled annotations to 100 licensed or trained professionals, ensuring clinical validity and reproducibility.
Adversarial Evaluation in Mental Health QA. Prior studies have highlighted risks of LLMs in mental health support, such as unsafe advice or discrepant responses based on inferred demographics (Lawrence et al., 2024; Imran et al., 2023; Gabriel et al., 2024). However, most findings come from post hoc analyses of deployed systems (Elyoseph & Levkovich, 2023), which reveal problems retrospectively rather than prospectively stress-testing models. A few recent works adapt red-teaming to mental health (Grabb et al., 2024; Schoene & Canca, 2025), but their failure modes are predefined and literature-driven, limiting coverage of the diverse failure patterns that emerge in practice. In contrast, our adversarial dataset is empirically grounded: 10 licensed mental health professionals authored 120 prompts from failures observed in CounselBench-Eval, creating an expert-crafted benchmark for prospectively eliciting clinically meaningful vulnerabilities.
@@ 3 Overview of CounselBench
We construct CounselBench in collaboration with 100 mental health professionals. It is a two-part benchmark for evaluating LLMs in realistic, mental health QA scenarios (Figure 1):
  • COUNSELBENCH-Eval: a benchmark of 2,000 expert evaluations of LLMs and human responses to real patient queries, with numeric scores, span-level annotations, and rationales.
@@
Figure 1: Overview of CounselBench benchmark. COUNSELBENCH-EVAL (left) includes expert evaluation of LLMs and human responses to real counseling questions. COUNSELBENCH-ADV (right) includes adversarial questions authored by clinicians to target identified LLM failure modes. See Appendix B for license/degree types and specialization areas.
Research Pipeline. We began by sampling 100 real patient questions from the public forum CounselChat (Bertagnolli, 2020), spanning 20 common mental health topics (§4.1). Each question was answered by GPT-4, Gemini, LLaMA-3, and a human therapist (§4.2). To evaluate responses, we defined six evidence-based dimensions grounded in literature: overall quality, empathy, specificity, factual consistency, medical advice, and toxicity (§4.3). We then recruited 100 mental health professionals for evaluation (§4.4). Besides assigning ratings, annotators provided span-level annotations and written rationales, yielding rich insights into response quality and safety. These evaluations form the COUNSELBENCH-EVAL dataset. We also prompted nine LLMs to rate the same responses under the same rubric (§4.6), enabling systematic comparison between human and LLM judges. To probe safety directly, 10 experts authored 120 adversarial questions, creating the CounselBench-Adv dataset (§5), to trigger failure patterns identified in CounselBench-Eval.
4 CounselBench-EVAL: Expert Evaluation of LLMs for Open‑Ended Mental‑Health QA
@@
We develop COUNSELBENCH-EVAL to systematically assess LLM behavior on authentic mental health questions. This section describes the construction pipeline: real-world mental health QA sourcing (§4.1), response generation (§4.2), rubric design (§4.3), human annotation protocol (§4.4), results from expert scoring and analysis (§4.5), and experiments with LLM-based evaluators (§4.6).
4.1 Real-World Mental Health Question and Answering
We source mental health QA pairs from CounselChat, an online platform where licensed therapists publicly respond to anonymous patient questions (Bertagnolli, 2020). All posts were authored before 2022, ensuring that all human responses predate the widespread use of LLMs like ChatGPT.
@@ 4.1 Real-World Mental Health Question and Answering
Question Curation and Topic Coverage. To ensure both topical diversity and response quality, we selected the top 20 topics based on predefined criteria (Appendix  LABEL:appendix_question_filtering). From each topic, we sampled five questions, resulting in 100 questions in total. The topics were: (1) depression, (2) relationships, (3) anxiety, (4) family conflict, (5) parenting, (6) self esteem, (7) relationship dissolution, (8) behavioral change, (9) anger management, (10) trauma, (11) marriage, (12) domestic violence, (13) grief and loss, (14) social relationships, (15) workplace relationships, (16) legal regulatory, (17) substance abuse, (18) counseling fundamentals, (19) eating disorders, (20) professional ethics.
Human Answers. Although all answers on CounselChat were written by certified clinicians, forum contributions are informal and vary in quality. To reduce this variability and focus on consistently well-regarded content, we selected the top-voted answer for each question (Appendix  LABEL:appendix_human_responses). Upvotes are a standard, publicly observable indicator of perceived helpfulness and relevance in peer-support forums, and have been widely used in prior work to capture perceived helpfulness(Jiang et al., 2020; Huang et al., 2025), social support (De Choudhury & De, 2014), engagement (Gao et al., 2020), community reception (Alghamdi et al., 2025), and narrative quality (Wang et al., 2017).
@@ 4.2 Generating Counseling Responses from LLMs
Pilot Study. We first piloted domain-specific models such as MentalLLaMA (Yang et al., 2023) and Meditron (Chen et al., 2023b). As shown in Appendix LABEL:sec:domain-specific-experiments, these models performed poorly on open-ended mental health QA, likely due to narrow fine-tuning or lack of instruction tuning. We therefore focused on general-purpose LLMs that represent current deployment paradigms
Model Selection. To ensure fair comparison and reduce annotator fatigue, we limited our evaluation to three LLMs per question. Each mental health professional blindly annotated four answers side by side (three LLM-generated and one human-written), enabling consistent evaluation while keeping the task cognitively manageable. The selected models are: (1) GPT-4-0613 (via OpenAI API), a widely adopted commercial model, (2) LLaMA-3.3-70B-Instruct (via Hugging Face), a leading open-source, instruction-tuned model from Meta, (3) Gemini-1.5-Pro (via Google Generative AI SDK), which powers responses in Google Search and Assistant. Together, they represent major deployment paradigms: commercial APIs, open-source checkpoints, and search-integrated platforms. Each model is prompted with the original patient question and asked to respond. Full prompt templates and generation details are provided in Appendix D.1.
@@ 4.3 Evaluation Rubric and Paradigms
We developed a multi-dimensional evaluation rubric based on clinical psychology literature and expert consultation. All metrics were rated using 5-point Likert scales ($1=$ the most negative; $5=$ the most positive) unless specified. Appendix A shows the rating interface.
@@
Empathy measures whether the response demonstrates emotional attunement, compassion, or validation. Empathy is a central component of therapeutic modalities (e.g., person-centered therapy (Jani et al., 2012)), and is strongly linked to therapeutic alliance, a key predictor of positive therapy outcomes (Ackerman & Hilsenroth, 2003).
Specificity evaluates how well the response is tailored to the user’s particular context, rather than offering generic or overly broad advice. Personalized replies help users feel heard and understood (Elkin et al., 2014; Stiles & Horvath, 2017; Kramer & Stiles, 2015), and contextual relevance is critical to effective support (Hitchcock et al., 2024; Stiles & Horvath, 2017; Kramer & Stiles, 2015).
Medical Advice flags whether the response includes therapeutic or diagnostic guidance that should only be provided by licensed professionals. Such advice typically requires individualized assessment and clinical expertise (Frank et al., 2020; Lyon et al., 2011), and its timing and appropriateness must be carefully managed by trained practitioners (Hill et al., 2023; Prass et al., 2021). This dimension is treated as binary (Yes/No) with an additional “I am not sure” option because fine-grained distinctions in severity and subcategories are difficult to assess consistently for Medical Advice. We further ask annotators to provide the exact advice spans and written rationales, which enable post hoc analysis of different categories, and levels of severity. Higher rates of flagged content indicate safety concerns.
Factual Consistency checks whether the response aligns with accepted commonsense or clinical knowledge and avoids inaccurate or unsupported claims. Users often struggle to discern factual accuracy in online mental health content (Morahan-Martin & Anderson, 2000). LLMs are known to hallucinate false information (Sun et al., 2025), and even human counselors can produce factual errors in communication (Lobel, 2024). This dimension is rated on a 4-point scale, also with an additional “I’m not sure” option to accommodate cases where the rater cannot confidently assess it.
Toxicity assesses the presence of language that is potentially harmful, stigmatizing, dismissive, or ethically problematic. Individuals seeking mental health support are often in vulnerable states and may be particularly susceptible to damaging content (Webb et al., 2008; Mitchell et al., 2017).
Rating Rationale & Text Extraction To capture the context and reasoning behind human evaluations of counseling messages, we asked annotators (§4.4) to provide open-text explanations for their ratings on Overall Quality, Medical Advice, Factual Consistency, and Toxicity. Additionally, we asked them to identify and extract specific portions of each response that supported their ratings. Two researchers on the team checked and verified the quality of text responses after data collection. In general, human annotators provided detailed written feedback, with an median length of 576.5 written words (IQR: 339.8 - 866.5) per survey combining all rationales.